# https://www.intelligence.tuc.gr/gogs/

### v2

@@ Abstract
This paper provides a comprehensive review of mainly Graph Neural Networks, Deep Reinforcement Learning, and Probabilistic Topic Modeling methods with a focus on their potential incorporation in strategic multiagent settings. We draw interest in (i) Machine Learning methods currently utilized for uncovering unknown model structures adaptable to the task of strategic opponent modeling, and (ii) the integration of these methods with Game Theoretic concepts that avoid relying on assumptions often invalid in real-world scenarios, such as the Common Prior Assumption (CPA) and the Self-Interest Hypothesis (SIH). We analyze the ability to handle uncertainty and heterogeneity, two characteristics that are very common in real-world application cases, as well as scalability. As a potential answer to effectively modeling relationships and interactions in multiagent settings, we champion the use of Graph Neural Networks (GNN). Such approaches are designed to operate upon graph-structured data, and have been shown to be a very powerful tool for performing tasks such as node classification and link prediction. Past related work indicates that GNNs are useful in cases of multiagent policy learning, and is a good step for overcoming restricting assumptions such as the CPA and SIH. Next, we review the domain of Reinforcement Learning (RL), and in particular that of Multiagent Deep Reinforcement Learning (MADRL). Single-agent deep RL (DRL) has been widely used for decision making in demanding game settings. Its application in multiagent settings though is hindered due to, e.g., varying relationships between agents, and non-stationarity of the environment. Following, we describe existing relevant game theoretic solution concepts and consider properties such as fairness and stability. Our review comes complete with a note on the literature that utilizes PTM in domains other than that of document analysis and classification. The capability of PTM to estimate unknown underlying distributions can help with tackling heterogeneity and unknown agent beliefs. Finally, we identify certain open challenges—specifically, the need to (i) fit non-stationary environments, (ii) balance the degrees of stability and adaptation, (iii) tackle uncertainty and heterogeneity, (iv) guarantee scalability and solution tractability.
@@ 1 Introduction
In recent years, the integration of MAS technologies has led to significant advancements in solving complex real-world problems [demazeau2010trends, 10.1007/978-3-540-75144-1_12, wooldridge2009introduction, xie2017multi]. Multiagent Systems (MAS) have garnered increasing attention due to their applicability in various domains such as for microeconomics, for trading in online markets, and for optimizing automated decision-making in cooperative and non-cooperative settings. The complexity of these systems arises from the interactions between agents [lutzenberger2016multi], the uncertainty in the environment [Nguyen_Kumar_Lau_2017, Varakantham_Cheng_Gordon_Ahmed_2021], and the need for coordination to achieve common or even conflicting goals [wooldridge2009introduction]. Traditional approaches often struggle to handle the dynamic nature and non-linear interactions inherent in Multiagent (MA) settings.
@@
MAS involve multiple agents interacting with each other in dynamic and often uncertain environments, requiring sophisticated approaches to coordination, decision-making, and learning. In such cases, the capability to accurately model opponents—or other players in general—is a crucial component [10.1007/3-540-60923-7_18, NEURIPS2022_b528459c]. Agents in MAS often need to adapt their strategies based on the actions and behaviors of other agents, thus it is desirable to predict the likely actions of their counterparts by learning from past interactions, enabling them to make informed decisions and anticipate possible outcomes. MA environments are typically dynamic, with agents continuously updating their strategies and responding to changes in the environment. Opponent modeling provides a mechanism for agents to track and to tackle these changes, helping them maintain effectiveness in dynamic and uncertain settings [weyns2007environment]. Moreover, understanding the strengths and weaknesses of opponents is essential for devising effective strategies so that agents can identify patterns, exploit vulnerabilities, and capitalize on the mistakes or suboptimal decisions made by their counterparts, thereby gaining a competitive advantage.
@@
In scenarios where agents need to collaborate or negotiate with each other, opponent modeling facilitates better understanding of the intentions, preferences, and behaviors of other agents. This understanding enables improved communication, coordination, and negotiation strategies, leading to more efficient and effective collaboration among agents. In parallel, uncertainty is inherent in multiagent systems due to incomplete information, stochasticity, and partial observability. Opponent modeling helps reduce uncertainty by providing agents with probabilistic estimates of opponent behaviors, thereby enabling more robust decision-making in uncertain environments. Finally, by observing and modeling the behaviors of opponents over time, agents can learn from experience and improve their strategies iteratively. This means that agents should be capable of updating their models dynamically based on new observations, feedback, and interactions, leading to continuous improvement and adaptation [weyns2009self].
@@
To this end, a series of methods have been proposed that aim to solve the problem of learning opponent beliefs [nyarko2002experimental, pmlr-v80-raileanu18a, zhai2023dynamic] and other agents’ internal model structures [858456, nashed2022survey]. According to such paradigms agents may observe the other agents’ past actions and payoffs, but the underlying belief models are private and must be inferred in some way. To achieve this, usually opponent agents are considered as separate entities that are independent from the environment, and a number of ‘exploratory’ interactions are required so as to gather the data required to build the corresponding models. Roughly, this process comprises a method for collecting data, a learning algorithm, and the decision regarding the target resolution, i.e. to learn the type of each agent, to predict the next actions of individuals, or to estimate the collective strategy of a team of agents. However, assuming that entities are independent from the environment is realistic only in very simple settings.
@@
As a possible solution, we examine the adoption of Bayesian GNNs [4700287, zhang2019bayesian], which have demonstrated effectiveness in diverse domains such as molecular structure learning, protein function prediction, disease classification, and speech recognition. Notably, recent applications have extended to predicting outcomes in turn-based strategy games and executing specific real-time strategy games using multi-agent deep reinforcement learning algorithms. However, to our knowledge, GNNs have yet to be leveraged for modeling intricate or uncertain strategic interaction domains or aiding in equilibrium learning within such contexts.
@@
Now, a large portion of existing opponent modeling methods relies on single-agent reinforcement learning (RL) with the aim to either maximize individual utility or the overall social welfare. However, the problem setting when transiting from single agent to multiagent learning environments changes substantially, as the latter poses more challenges. Specifically, theoretical convergence guarantees are hindered due to non-stationarity, incomplete information, and the fact that different agent actions change the states of the environment in ways that are hard to estimate. Also, it is not certain that the Markov assumption continues to hold. Taking these into account, more types of multiagent learning methods have been applied, including unsupervised learning techniques, e.g. that employ PTM, and other methods that model the whole agent population policies, e.g. co-evolutionary policy learning techniques, swarm intelligence, or learn the interaction protocols in effect, e.g. via adaptive mechanism design [10.1007/978-3-030-01713-2_1].
@@
The particular challenges that arise in strategic MA settings have been a subject of study by past research. One important issue is the exponential growth of the computational complexity of classic RL algorithms when the state-action pairs come in large numbers. In MAS this is further worsened as the number of agents is an additional factor that adds to this exponential growth [Busoniu2010]. Another issue is that of the resolution of the different agent utility and payoffs, since these are typically correlated and their independent maximization is often not realizable. Moreover, the various agents learn simultaneously as they act in the environment leading to the problem of the best policy changing dynamically and not being a stationary point. Adding to this, the exploration-exploitation trade-off becomes even more difficult to obtain in such multiagent settings and it is more probable to destabilize the learning process [Busoniu2010].
@@
One can categorize multiagent reinforcement learning (MARL) into three types [tuyls2012multiagent]: multiplied learning, divided learning, and interactive learning. In the first case, the learning task is performed independently by each agent and the interactions among them do not affect the way they learn. The resulting strategies can be considered a generalization that can be applied in any setting. In the divided learning case, each agent is assigned particular tasks and the result can be combined to achieve a common goal. Contrary to the previous case, the strategies that each agent learns correspond to different areas or domains, and, if combined, can lead to social welfare maximization. Finally, interactive learning techniques can be considered more sophisticated than the other categories, since this type of learning aims to strike a balance between the two previous extremes, generalization and specialization. In this case the learning tasks are not known a priori, and agents need to explore and negotiate a successful learning path. Agents regulate the joint learning procedure and then synthesize the various perspectives and conflicts that the learning results may pose.
@@
Furthermore, DRL has emerged in the recent years as a powerful tool for decision making in strategic games, such as Go [mnih2015human, silver2017mastering]. DRL essentially employs the power of modern deep networks to perform function approximation in reinforcement learning environments [sutton2018reinforcement]. However, DRL methods have so far been used mainly in two- and not multi- player strategic games; and typically require a large number of learning experiences to reach a satisfactory level of performance.
@@
Overall, in MARL, agents take part in a stochastic game with unknown reward and/or transition models [littman1994markov, littman2001value, shapley1953stochastic], and must learn to (inter)act optimally in the presence of the other agents doing the same, and also considering the incompleteness of information, the vast sizes of state spaces, and the distribution of rewards. The agents try to maximize their expected discounted rewards; while their optimal behavior can be defined in terms of game theoretic solution concepts (such as some variant of the well-known Nash equilibrium) [shoham2007if]. However, note that demanding convergence of the learning process to some sort of equilibria should not necessarily be seen as a requirement or even as a “nice-to-have” property in real-world multiagent systems [stone2007multiagent]. In general, MARL should not be considered just as the multiagent extension of the reinforcement learning problem. Rather, in most if not all cases, a MARL method should be explicitly taking other agents into account; and, at the same time, one could be using tools provided by, e.g., game theory to reason about others and thus improve their own performance.
@@
Notwithstanding it being widely employed in modern MAS and AI research [chalkiadakis2022computational, shoham2008multiagent], modern algorithmic game theory (AGT) [roughgarden2010algorithmic] has been criticized for offering, more often than not, negative computational complexity results. Moreover, key game-theoretic solution concepts, such as the celebrated NE, have been criticized for (a) being incapable of characterizing the nature of most real-world equilibria situations as argued [daskalakis2008complexity]; and for (b) constituting a controversial “target” or “holy grail” for rational behavior in multiagent learning environments [chalkiadakis2003coordination, shoham2007if, stone2007multiagent]. These criticisms are based mainly on the high computational complexity of NE and related concepts, and to the often unrealistic full-rationality, or other simplifying assumptions adopted in GT. Indeed, even if one would be able to compute equilibria within a reasonable time-frame, the rationality of the equilibrium-prescribed behaviour would often be questionable, given the assumptions usually present in existing equilibrium notions. Therefore, our view is that more steps need to be taken to remove restrictive GT assumptions and practices, and integrate GT further with modern AI and MAS research, to practical benefits.
@@
One specific assumption that is central to the usually studied equilibrium notions is the CPA—namely, the assumption that the beliefs of the various agents are posteriors, obtained by conditioning a common prior on private information received by the individuals. The common prior assumption is a substantive assumption, and alongside certain common knowledge assumptions, it plays an important role in game theory and economics: specifically, it is central to the definition of Bayesian (or “Bayes-Nash”) equilibria [shoham2008multiagent], and is used to restrict attention to correlated equilibria, and (when beliefs are uncorrelated) to NE. As such, the CPA has provided the main decision-theoretic justification of such concepts (see, e.g., [aumann1987correlated]). Convenient as this assumption might be, prior information is not, in fact, common in most real-world environments. Moreover, there exist problem settings where explicitly modeling prior beliefs as heterogeneous can improve our understanding and tackling of such problems [morris1995common].
@@
Nevertheless, arguments in favor of the CPA’s plausibility exist, such as that if people have always been fed the same information, then it is not unreasonable to assume a common prior [aumann1987correlated]. However, the CPA’s implication that there exists one ex ante, prior stage during which all agents share the same beliefs regarding each other’s beliefs, including beliefs regarding their own actions (and regarding how these beliefs are going to be revised), is, arguably, highly improbable [gul1998comment, morris1995common]. Moreover, the assumption of a commonly shared prior during some prior stage of deliberations, runs counter to the Bayesian view of statistics, and of interpreting information and beliefs [jaynes2003probability], since it amounts to asserting that at some moment in time everyone must have had identical beliefs.
@@
To the best of our knowledge, all criticisms against the CPA and all attempts to replace its problematic conceptual elements, have so far been provided from a mainly philosophical/theoretical viewpoint. To date, there has been no convincing attempt to override CPA in practice, through explicitly considering a truly dynamic stage through which agents form their initial beliefs. Moreover, one of the most compelling arguments in favour of the CPA is that it considerably simplifies analysis [morris1995common]. However, we argue this need not be the case: the use of modern Bayesian techniques can ease the analysis, even under heterogeneous priors.
@@
Moreover, another widely used assumption which has been questioned is the SIH according to which agents’ objective is the maximization of their personal payoff. Namely, experiments conducted primarily in the 1980s and 1990s demonstrated that people are not in general self-interested and, as a result, many scientists proposed models which attempt to explain this phenomenon by introducing in agents’ decision- making processes the notions of reciprocity, fairness and altruism to name a few [fehr2006economics, sobel2005interdependent]. Nevertheless, economists have tried to provide a justification for deviations from solution concepts such as the NE, and never tried to adopt a bottom-up approach, e.g. from agent characteristics to agent interactions. We anticipate that by using machine learning techniques in order to reveal the structure of an agent’s thinking and reasoning this can be achieved.
@@
Next, in Section 3 we analyze the most recent and effective RL algorithms, with a focus also on MA applications. DRL techniques have shown remarkable success in learning optimal policies for individual agents in complex environments. When applied to multiagent settings, DRL algorithms enable agents to learn strategies through interaction with the environment and other agents.

### v3

@@ Abstract
This paper provides a comprehensive review of mainly Graph Neural Networks, Deep Reinforcement Learning, and Probabilistic Topic Modeling methods with a focus on their potential incorporation in strategic multiagent settings. We draw interest in (i) Machine Learning methods currently utilized for uncovering unknown model structures adaptable to the task of strategic opponent modeling, and (ii) the integration of these methods with Game Theoretic concepts that avoid relying on assumptions often invalid in real-world scenarios, such as the Common Prior Assumption (CPA) and the Self-Interest Hypothesis (SIH). We analyze the ability to handle uncertainty and heterogeneity, two characteristics that are very common in real-world application cases, as well as scalability. As a potential answer to effectively modeling relationships and interactions in multiagent settings, we champion the use of Graph Neural Networks (GNN). Such approaches are designed to operate upon graph-structured data, and have been shown to be a very powerful tool for performing tasks such as node classification and link prediction. Past related work indicates that GNNs are useful in cases of multiagent policy learning, and is a good step for overcoming restricting assumptions such as the CPA and SIH. Next, we review the domain of reinforcement learning (RL), and in particular that of multiagent deep reinforcement learning (MADRL). Single-agent deep RL (DRL) has been widely used for decision making in demanding game settings. Its application in multiagent settings though is hindered due to, e.g., varying relationships between agents, and non-stationarity of the environment. We describe existing relevant game theoretic solution concepts, and consider properties such as fairness and stability. Our review comes complete with a note on the literature that utilizes probabilistic topic modeling (PTM) in domains other than that of document analysis and classification. The ability of probabilistic topic models to estimate unknown underlying distributions can help with tackling heterogeneity and unknown agent beliefs. Finally, we identify certain open challenges—specifically, the need to (i) fit non-stationary environments, (ii) balance the degrees of stability and adaptation, (iii) tackle uncertainty and heterogeneity, (iv) guarantee scalability and solution tractability.
@@ 1 Introduction
In recent years, multiagent systems (MAS) [149] have gathered increasing attention due to their applicability in various domains such as for microeconomics, for trading in online markets, for optimizing automated decision-making in cooperative and non-cooperative settings, and in general for leading to significant advancements towards the solution of complex real-world problems [35, 48, 151]. The complexity of these systems arises from the interactions between agents [84], the uncertainty in the environment [97, 139], and the need for coordination to achieve common or even conflicting goals [149]. Traditional approaches often struggle to handle the dynamic nature and non-linear interactions inherent in multiagent (MA) settings.
@@
MAS involve multiple agents interacting with each other in dynamic and often uncertain environments, requiring sophisticated approaches to coordination, decision-making, and learning. In such cases, the capability to accurately model opponents—or other players in general—is a crucial component [20, 158]. Agents in MAS often need to adapt their strategies based on the actions and behaviors of other agents, thus it is desirable to predict the likely actions of their counterparts by learning from past interactions, enabling them to make informed decisions and anticipate possible outcomes. MA environments are typically dynamic, with agents continuously updating their strategies and responding to changes in the environment. Opponent modeling [2] provides a mechanism for agents to track and to tackle these changes, helping them maintain effectiveness in dynamic and uncertain settings. Moreover, understanding the strengths and weaknesses of opponents is essential for devising effective strategies so that agents can identify patterns, exploit vulnerabilities, and capitalize on the mistakes or suboptimal decisions made by their counterparts, thereby gaining a competitive advantage.
@@
In scenarios where agents need to collaborate or negotiate with each other, opponent modeling facilitates better understanding of the intentions, preferences, and behaviors of other agents. This understanding enables improved communication, coordination, and negotiation strategies, leading to more efficient and effective collaboration among agents. In parallel, uncertainty is inherent in multiagent systems due to incomplete information, stochasticity, and partial observability. Opponent modeling helps reduce uncertainty by providing agents with probabilistic estimates of opponent behaviors, thereby enabling more robust decision-making in uncertain environments. Finally, by observing and modeling the behaviors of opponents over time, agents can learn from experience and improve their strategies iteratively. This means that agents should be capable of updating their models dynamically based on new observations, feedback, and interactions, leading to continuous improvement and adaptation [147].
@@
To this end, a series of methods have been proposed that aim to solve the problem of learning the beliefs of one’s opponent(s) [98, 110, 159], as well as the internal model structures of other agents [95, 96]. According to such paradigms agents may observe the other agents’ past actions and payoffs (or, interestingly, observations and actions of the modelled agents may be replicated relying only on local information of the controlled agent [105]); but the underlying belief models are private and must be inferred in some way. To achieve this, usually opponent agents are considered as separate entities that are independent from the environment, and a number of “exploratory” interactions are required so as to gather the data required to build the corresponding models. Roughly, this process comprises a method for collecting data, a learning algorithm, and the decision regarding the target resolution, i.e. to learn the type of each agent, to predict the next actions of individuals, or to estimate the collective strategy of a team of agents. However, assuming that entities are independent from the environment is realistic only in very simple settings.
@@
As a possible solution, we examine the adoption of Graph Neural Networks (GNNs) [115], which have demonstrated effectiveness in diverse domains such as molecular structure learning, protein function prediction, disease classification, and speech recognition. Notably, recent applications have extended to predicting outcomes in turn-based strategy games and executing specific real-time strategy games using multi-agent deep reinforcement learning algorithms. However, to our knowledge, GNNs have yet to be leveraged for modeling intricate or uncertain strategic interaction domains or aiding in equilibrium learning within such contexts.
@@
Now, a large portion of existing opponent modeling methods relies on single-agent reinforcement learning (RL) with the aim to either maximize individual utility or the overall social welfare. However, the problem setting when transiting from single agent to multiagent learning environments changes substantially, as the latter poses more challenges. Specifically, theoretical convergence guarantees are hindered due to non-stationarity, incomplete information, and the fact that different agent actions change the states of the environment in ways that are hard to estimate. Also, it is not certain that the Markov assumption continues to hold. Taking these into account, more types of multiagent learning methods have been applied, including unsupervised learning techniques, e.g. that employ PTM, and other methods that model the whole agent population policies, e.g. co-evolutionary policy learning techniques, swarm intelligence, or learn the interaction protocols in effect, e.g. via adaptive mechanism design [137].
@@
The particular challenges that arise in strategic MA settings have been a subject of study by past research. One important issue is the exponential growth of the computational complexity of classic RL algorithms when the state-action pairs come in large numbers. In MAS this is further worsened as the number of agents is an additional factor that adds to this exponential growth [18]. Another issue is that of the resolution of the different agent utility and payoffs, since these are typically correlated and their independent maximization is often not realizable. Moreover, the various agents learn simultaneously as they act in the environment leading to the problem of the best policy changing dynamically and not being a stationary point. Adding to this, the exploration-exploitation trade-off becomes even more difficult to obtain in such multiagent settings and it is more probable to destabilize the learning process [18].
@@
One can categorize multiagent reinforcement learning (MARL) into three types [138]: multiplied learning, divided learning, and interactive learning. In the first case, the learning task is performed independently by each agent and the interactions among them do not affect the way they learn. The resulting strategies can be considered a generalization that can be applied in any setting. In the divided learning case, each agent is assigned particular tasks and the result can be combined to achieve a common goal. Contrary to the previous case, the strategies that each agent learns correspond to different areas or domains, and, if combined, can lead to social welfare maximization. Finally, interactive learning techniques can be considered more sophisticated than the other categories, since this type of learning aims to strike a balance between the two previous extremes, generalization and specialization. In this case the learning tasks are not known a priori, and agents need to explore and negotiate a successful learning path. Agents regulate the joint learning procedure and then synthesize the various perspectives and conflicts that the learning results may pose.
@@
Furthermore, deep reinforcement learning (DRL) has emerged in the recent years as a powerful tool for decision making in strategic games, such as Go [90, 127]. DRL essentially employs the power of modern deep networks to perform function approximation in reinforcement learning environments [134]. However, DRL methods have so far been used mainly in two- and not multi- player strategic games; and typically require a large number of learning experiences to reach a satisfactory level of performance.
@@
Overall, in MARL, agents take part in a stochastic game with unknown reward and/or transition models [77, 78, 121], and must learn to (inter)act optimally in the presence of the other agents doing the same, and also considering the incompleteness of information, the vast sizes of state spaces, and the distribution of rewards. The agents try to maximize their expected discounted rewards; while their optimal behavior can be defined in terms of game theoretic solution concepts (such as some variant of the well-known Nash equilibrium) [125]. However, note that demanding convergence of the learning process to some sort of equilibria should not necessarily be seen as a requirement or even as a “nice-to-have” property in real-world multiagent systems [131]. In general, MARL should not be considered just as the multiagent extension of the reinforcement learning problem. Rather, in most if not all cases, a MARL method should be explicitly taking other agents into account; and, at the same time, one could be using tools provided by, e.g., game theory to reason about others and thus improve their own performance.
@@
Notwithstanding it being widely employed in modern MAS and AI research [24, 124], modern algorithmic game theory (AGT) [111] has been criticized for offering, more often than not, negative computational complexity results. Moreover, key game-theoretic solution concepts, such as the celebrated NE, have been criticized for (a) being incapable of characterizing the nature of most real-world equilibria situations as argued [32]; and for (b) constituting a controversial “target” or “holy grail” for rational behavior in multiagent learning environments [22, 125, 131]. These criticisms are based mainly on the high computational complexity of NE and related concepts, and to the often unrealistic full-rationality, or other simplifying assumptions adopted in GT. Indeed, even if one would be able to compute equilibria within a reasonable time-frame, the rationality of the equilibrium-prescribed behaviour would often be questionable, given the assumptions usually present in existing equilibrium notions. Therefore, our view is that more steps need to be taken to remove restrictive GT assumptions and practices, and integrate GT further with modern AI and MAS research, to practical benefits.
@@
One specific assumption that is central to the usually studied equilibrium notions is the CPA—namely, the assumption that the beliefs of the various agents are posteriors, obtained by conditioning a common prior on private information received by the individuals. The common prior assumption is a substantive assumption, and alongside certain common knowledge assumptions, it plays an important role in game theory and economics: specifically, it is central to the definition of Bayesian (or “Bayes-Nash”) equilibria [124], and is used to restrict attention to correlated equilibria, and (when beliefs are uncorrelated) to NE. As such, the CPA has provided the main decision-theoretic justification of such concepts (see, e.g., [7]). Convenient as this assumption might be, prior information is not, in fact, common in most real-world environments. Moreover, there exist problem settings where explicitly modeling prior beliefs as heterogeneous can improve our understanding and tackling of such problems [93].
@@
Nevertheless, arguments in favor of the CPA’s plausibility exist, such as that if people have always been fed the same information, then it is not unreasonable to assume a common prior [7]. However, the CPA’s implication that there exists one ex ante, prior stage during which all agents share the same beliefs regarding each other’s beliefs, including beliefs regarding their own actions (and regarding how these beliefs are going to be revised), is, arguably, highly improbable [51, 93]. Moreover, the assumption of a commonly shared prior during some prior stage of deliberations, runs counter to the Bayesian view of statistics, and of interpreting information and beliefs [61], since it amounts to asserting that at some moment in time everyone must have had identical beliefs.
@@
To the best of our knowledge, all criticisms against the CPA and all attempts to replace its problematic conceptual elements, have so far been provided from a mainly philosophical/theoretical viewpoint. To date, there has been no convincing attempt to override CPA in practice, through explicitly considering a truly dynamic stage through which agents form their initial beliefs. Moreover, one of the most compelling arguments in favour of the CPA is that it considerably simplifies analysis [93]. However, we argue this need not be the case: the use of modern Bayesian techniques can ease the analysis, even under heterogeneous priors.
@@
Moreover, another widely used assumption which has been questioned is the SIH according to which agents’ objective is the maximization of their personal payoff. Namely, experiments conducted primarily in the 1980s and 1990s demonstrated that people are not in general self-interested and, as a result, many scientists proposed models which attempt to explain this phenomenon by introducing in agents’ decision- making processes the notions of reciprocity, fairness and altruism to name a few [40, 129]. Nevertheless, economists have tried to provide a justification for deviations from solution concepts such as the NE, and never tried to adopt a bottom-up approach, e.g. from agent characteristics to agent interactions. We anticipate that by using machine learning techniques in order to reveal the structure of an agent’s thinking and reasoning this can be achieved.
@@
Next, in Section 3 we analyze the most recent and effective RL algorithms, with a focus also on MA applications. Deep reinforcement learning techniques have shown remarkable success in learning optimal policies for individual agents in complex environments. When applied to multiagent settings, DRL algorithms enable agents to learn strategies through interaction with the environment and other agents.